OpenStudy (anonymous):

find the slope intercept form for the line satisfying the following conditions x intercept 3, y intercept 2/3

OpenStudy (anonymous):

so you have two points:(3,0) and (0,2/3). Can you continue question after that?

OpenStudy (anonymous):

Are you able to continue answer after this?

OpenStudy (anonymous):

i am trying it right now

OpenStudy (anonymous):

I am still a little stumped on this one. the 2/3 is throwing me off

OpenStudy (anonymous):

ok.the equation you want is of form y=mx+c m=(y2-y1)/(x2-x1) m=-2/9 we c is a y-intercept and is equal to 2/3 so y=(-2/9)x+2/3